# The impact of finite span and wing-tip vortices on a turbulent NACA0012 wing

### v3

@@ 2.3 Computational grids
RWT-0 $952\times 10^{6}$ $16\times 10^{-6}$ (11.4, 0.70, 6.9) $(11,1.7,2.2)\times 10^{-4}$ (5.5, 2.6, 1.7)
RWT-5 $1.56\times 10^{9}$ $8.5\times 10^{-6}$ (10.6, 0.75, 6.0) $(11,1.7,2.2)\times 10^{-4}$ (5.7, 2.8, 1.7)
RWT-10 $2.16\times 10^{9}$ $6.3\times 10^{-6}$ (10.6, 0.80, 5.4) $(11,1.7,2.2)\times 10^{-4}$ (5.6, 3.0, 1.8)
@@ 2.4 Flow transients and statistical averaging
After the transient period, turbulence statistics are collected over a period of $5\mathcal{T}_{\rm conv}$ or above in the periodic cases (P-0, P-2, P-5, P-10), around $8\mathcal{T}_{\rm conv}$ for RWT-0 and RWT-10, and for a longer period of $23.2\mathcal{T}_{\rm conv}$ in RWT-5. Table 2 summarizes the averaging time $t_{\rm avg}$ used in each simulation in terms of the convective time unit. A conversion ratio is provided to relate $\mathcal{T}_{\rm conv}$ to other time scales of the flow, including: the boundary layer eddy-turnover time $\mathcal{T}_{\rm ETT}=\delta_{99}/u_{\tau}$, where $\delta_{99}$ is the $99\%$ boundary layer thickness and $u_{\tau}=\sqrt{\tau_{w}/\rho}$ is the friction velocity; the shear-layer (wake) time scale $\mathcal{T}_{\rm shear}=\delta_{\rm 0.5,shear}/U_{\infty}$, where $\delta_{\rm 0.5,shear}$ is a measure of the wake thickness (see Pope, 2000); the vortex-rotation time scale $\mathcal{T}_{\rm vort}=\pi d_{\rm vort}/u_{\theta,{\rm max}}$, where $u_{\theta,{\rm max}}$ is the maximum azimuthal velocity around the vortex core and $d_{\rm vort}$ is the vortex diameter measured as the distance between the two peaks in the azimuthal velocity around the core. The RWT-0 (which is symmetric along $y$ and can be averaged in that direction) and RWT-5 have similar effective values of $t_{\rm avg}/\mathcal{T}_{\rm ETT,ss}$ (which is the most relevant time scale for the quantities studied here), while RWT-10 has a shorter integration time due to computational-cost constraints.
@@
In each case, the full statistical record is divided into four batches of equal size, which are used to estimate the uncertainty in solution statistics by computing the confidence intervals of each quantity of interest using the non-overlapping batch method (cf. Conway, 1963). Given the higher sensitivity of the Reynolds stresses, particularly for finite-span wings, their approximate error bars are included in the comparisons in section 4.3 and appendix C. With the use of ensemble averaging along the span in periodic cases and the equivalent filtering in the wing-tip cases, the averaging times used in this work are deemed sufficient for the discussions here.
@@ 4.1 The impact of the effective angle of attack
Figure 6: The $99\%$ boundary layer thickness $\delta_{99}$ (a, b) and the Clauser pressure-gradient parameter $\beta_{x_{\rm BL}}$ (c, d) plotted at $z^{\prime}/c=0.1$ (close to the root) on the suction side (a, c) and pressure side (b, d) of the wings of this study. Different colors illustrate angles of attack of $\alpha=0^{\circ}$, $2^{\circ}$, $5^{\circ}$, and $10^{\circ}$, with darker colors corresponding to higher angles of attack and matched colors for matched $\alpha$. The finite-span wings are plotted in solid lines, while periodic airfoils are shown by the dotted lines. The observed differences are mainly attributed to the impact of the reduced effective angle of attack $\alpha_{\rm eff}$ (figure 1) on the boundary layers.
There is a large difference between the boundary layers formed on the finite-span wings compared to the periodic ones at the same angle of attack. Namely, the boundary layers on the suction sides are thinner and encounter lower adverse pressure gradients (lower $\beta_{x_{\rm BL}}$ values), while they are thicker on the pressure side and encounter a larger Clauser parameter (interesting to note that $\beta_{x_{\rm BL}}>0$ over the majority of the pressure side as well). Furthermore, the discrepancies are larger for higher angles of attack and nearly vanish for $\alpha=0^{\circ}$. These are all consequences of the lower effective angle of attack $\alpha_{\rm eff}$ of the finite-span wings compared to their geometric angle of attack $\alpha$ ( a well-known inviscid effect explained in figure 1). This is further demonstrated by figure 7 which shows the section-wise pressure component of the lift and drag coefficients of all cases defined as,
@@
Figure 7: The section-wise (a) pressure lift and (b) pressure drag coefficients along the span for all the geometries of this study. Line styles and colors are the same as those in figure 6. Refer to figure 1 (b) and its caption for an explanation of the increased drag coefficient of the finite-span wings.
@@ 4.3 Additional departures from infinite-span wings
mm 0mm]{a5rwt_ss_RSplus_z1_allX.eps}
[…]
mm 0mm]{a5rwt_ss_RSplus_z4_allX.eps}
[…]
mm 0mm]{a5rwt_ss_RSplus_x0.7_allZ.eps}
[…]
mm 0mm]{a5rwt_ss_RSplus_x0.9_allZ.eps}
[…]
@@ Appendix C More details on the comparison of boundary layer quantities for finite- and infinite-span wings
$x^{\prime}=0.6$ $\frac{(241,1.32,0.05)}{(249,1.26)}$ $\frac{(238,1.27,0.09)}{(249,1.24)}$ $\frac{(237,1.16,0.12)}{(244,1.13)}$ $\frac{(244,1.00,0.05)}{(236,1.04)}$
$x^{\prime}=0.7$ $\frac{(265,1.70,0.03)}{(266,1.70)}$ $\frac{(263,1.63,0.04)}{(265,1.63)}$ $\frac{(264,1.48,0.01)}{(261,1.48)}$ $\frac{(272,1.22,-0.16)}{(254,1.33)}$
$x^{\prime}=0.8$ $\frac{(283,2.48,0.00)}{(278,2.48)}$ $\frac{(284,2.38,-0.03)}{(277,2.37)}$ $\frac{(287,2.20,-0.19)}{(277,2.21)}$ $\frac{(294,1.75,-0.54)}{(272,1.83)}$
$x^{\prime}=0.9$ $\frac{(293,4.15,-0.06)}{(282,4.17)}$ $\frac{(294,4.09,-0.18)}{(282,4.08)}$ $\frac{(293,3.97,-0.46)}{(282,3.95)}$ $\frac{(296,3.29,-1.19)}{(284,3.28)}$

### v4

@@ 2.3 Computational grids
RWT-0 $952\times 10^{6}$ $16\times 10^{-6}$ (11.4, 0.83, 6.4) $(11,1.7,2.2)\times 10^{-4}$ (5.5, 2.6, 1.7)
RWT-5 $1.56\times 10^{9}$ $8.5\times 10^{-6}$ (10.6, 0.77, 5.9) $(11,1.7,2.2)\times 10^{-4}$ (5.6, 2.8, 1.8)
RWT-10 $2.16\times 10^{9}$ $6.3\times 10^{-6}$ (9.9, 0.72, 5.5) $(11,1.7,2.2)\times 10^{-4}$ (5.6, 3.0, 1.9)
@@ 2.4 Flow transients and statistical averaging
After the transient period, turbulence statistics are collected over a period of $5\mathcal{T}_{\rm conv}$ or above in the periodic cases (P-0, P-2, P-5, P-10), around $8\mathcal{T}_{\rm conv}$ for RWT-0, around $14\mathcal{T}_{\rm conv}$ for RWT-10, and for a longer period of around $23\mathcal{T}_{\rm conv}$ in RWT-5. Table 2 summarizes the averaging time $t_{\rm avg}$ used in each simulation in terms of the convective time unit. A conversion ratio is provided to relate $\mathcal{T}_{\rm conv}$ to other time scales of the flow, including: the boundary layer eddy-turnover time $\mathcal{T}_{\rm ETT}=\delta_{99}/u_{\tau}$, where $\delta_{99}$ is the $99\%$ boundary layer thickness and $u_{\tau}=\sqrt{\tau_{w}/\rho}$ is the friction velocity; the shear-layer (wake) time scale $\mathcal{T}_{\rm shear}=\delta_{\rm 0.5,shear}/U_{\infty}$, where $\delta_{\rm 0.5,shear}$ is a measure of the wake thickness (see Pope, 2000); the vortex-rotation time scale $\mathcal{T}_{\rm vort}=\pi d_{\rm vort}/u_{\theta,{\rm max}}$, where $u_{\theta,{\rm max}}$ is the maximum azimuthal velocity around the vortex core and $d_{\rm vort}$ is the vortex diameter measured as the distance between the two peaks in the azimuthal velocity around the core. The RWT-0 (which is symmetric around the $y=0$ plane and can be averaged in that direction) and RWT-5 have similar effective values of $t_{\rm avg}/\mathcal{T}_{\rm ETT,ss}$ (which is the most relevant time scale for the quantities studied here), while RWT-10 has a shorter integration time due to computational-cost constraints.
@@
In each case, the full statistical record is divided into four or more batches of equal size, which are used to estimate the uncertainty in solution statistics by computing the confidence intervals of each quantity of interest using the non-overlapping batch method (cf. Conway, 1963). Given the higher sensitivity of the Reynolds stresses, particularly for finite-span wings, their approximate error bars are included in the comparisons in section 4.3 and appendix C. With the use of ensemble averaging along the span in periodic cases and the equivalent filtering in the wing-tip cases, the averaging times used in this work are deemed sufficient for the discussions here.
@@ 4.1 The impact of the effective angle of attack
Figure 6: The $99\%$ boundary layer thickness $\delta_{99}$ (a, b) and the Clauser pressure-gradient parameter $\beta_{x_{\rm BL}}$ (c, d) plotted at $z^{\prime}/c=0.1$ (close to the root) on the suction side (a, c) and pressure side (b, d) of the wings of this study. The finite-span wings are plotted in solid lines, while periodic airfoils are shown by the dotted lines. Colors, from light to dark, correspond to angles of attack of $\alpha=0^{\circ}$, $2^{\circ}$, $5^{\circ}$, and $10^{\circ}$, respectively, for infinite-span wings. The three finite-span wings are shown with matched colors to the infinite-span ones at the same $\alpha$. The observed differences are mainly attributed to the impact of the reduced effective angle of attack $\alpha_{\rm eff}$ (an inviscid effect caused by the induced downwash of the wing-tip vortex; see figure 1) on the boundary layers.
There is a large difference between the boundary layers formed on the finite-span wings compared to the periodic ones at the same angle of attack. Namely, the boundary layers on the suction sides are thinner and encounter lower adverse pressure gradients (lower $\beta_{x_{\rm BL}}$ values), while they are thicker on the pressure side and encounter a larger Clauser parameter (interesting to note that $\beta_{x_{\rm BL}}>0$ over the majority of the pressure side as well). Furthermore, the discrepancies are larger for higher angles of attack and nearly vanish for $\alpha=0^{\circ}$. These are all consequences of the lower effective angle of attack $\alpha_{\rm eff}$ of the finite-span wings compared to their geometric angle of attack $\alpha$ (a well-known inviscid effect caused by the induced downwash of the wing-tip vortex, explained in figure 1). This is further demonstrated by figure 7 which shows the section-wise pressure component of the lift and drag coefficients of all cases defined as,
@@
Figure 7: The section-wise (a) pressure lift and (b) pressure drag coefficients along the span for all the geometries of this study. Line styles and colors are the same as those in figure 6. The additional thin solid lines in panel (a) are from inviscid simulations of finite-span wings for the same configurations as RWT-0, RWT-5, and RWT-10. The nearly identical variation of pressure lift coefficient (approximately proportional to $\alpha_{\rm eff}$) emphasizes the inviscid nature of the phenomenon. Refer to figure 1 (b) and its caption for a schematic explanation of the induced drag which explains the increased drag coefficient of the finite-span wings.
@@
Figure 7 (a) also compares the turbulent finite-span wings, RWT-0, RWT-5, and RWT-10, with inviscid (incompressible Euler) simulations of the same setups (without the tripping) performed using the open-source solver SU2 (Economon et al., 2016). The nearly identical variation of the lift (and effective angle of attack) over a large portion of the span emphasizes the inviscid nature of the phenomenon. The larger observed differences near the tip should be mainly attributed to the difference in vortex formation and the resulting change in both the vertical and spanwise location of the vortex in inviscid flows. These differences are discussed briefly in appendix B. Additional viscous effects are present near the tip and are expected to contribute to the observed differences in that region; however, they will not be discussed here.
@@ Appendix B The wing-tip region
An important distinction should be made between the vortex formation described here and that of an inviscid flow. In particular, in the absence of viscosity and no-slip boundary conditions at the wall the wing-tip vortex only separates farther downstream near the tip of the trailing-edge where the radius of curvature approaches zero (cf. Rizzi & Eriksson, 1984). The change in the location of the vortex leads to variations in the induced downwash on the wing. As was observed in figure 7 such variations are rather small away from the vortex core and become significant only closer to the wing tip. Additionally, the formation of secondary and tertiary vortices are not observed in inviscid flows.
@@ Appendix C More details on the comparison of boundary layer quantities for finite- and infinite-span wings
$x^{\prime}=0.6$ ${\color[rgb]{0,0,0}\frac{(240,1.33,0.05)}{(249,1.26)}}$ ${\color[rgb]{0,0,0}\frac{(239,1.27,0.09)}{(248,1.24)}}$ ${\color[rgb]{0,0,0}\frac{(237,1.16,0.12)}{(244,1.13)}}$ ${\color[rgb]{0,0,0}\frac{(245,0.99,0.05)}{(237,1.05)}}$
$x^{\prime}=0.7$ ${\color[rgb]{0,0,0}\frac{(263,1.72,0.03)}{(266,1.70)}}$ ${\color[rgb]{0,0,0}\frac{(264,1.63,0.04)}{(265,1.64)}}$ ${\color[rgb]{0,0,0}\frac{(265,1.47,0.00)}{(261,1.48)}}$ ${\color[rgb]{0,0,0}\frac{(273,1.21,-0.17)}{(254,1.33)}}$
$x^{\prime}=0.8$ ${\color[rgb]{0,0,0}\frac{(283,2.48,0.01)}{(278,2.48)}}$ ${\color[rgb]{0,0,0}\frac{(284,2.40,-0.03)}{(278,2.41)}}$ ${\color[rgb]{0,0,0}\frac{(289,2.19,-0.19)}{(277,2.19)}}$ ${\color[rgb]{0,0,0}\frac{(294,1.76,-0.55)}{(272,1.83)}}$
$x^{\prime}=0.9$ ${\color[rgb]{0,0,0}\frac{(292,4.20,-0.04)}{(283,4.22)}}$ ${\color[rgb]{0,0,0}\frac{(292,4.17,-0.18)}{(282,4.17)}}$ ${\color[rgb]{0,0,0}\frac{(298,3.85,-0.46)}{(282,3.87)}}$ ${\color[rgb]{0,0,0}\frac{(297,3.31,-1.19)}{(284,3.31)}}$